The area of a rectangle is 30cm squared
its lenghth and width are whole numbers
Find the smallest possible permimeter of this rectangle
This question is 4 marks

Answer by Boyen(12)

perimeter = (l+w)*2
area = w*l
Out of this we learn that you want to look for 2 divisors that counted together are as low as possible
all divisors of 30 :
1*30 , 2*15 , 3*10, 5*6, 6*5,
5*6 is the answer because 5+6 = only 11
so perimeter = 11*2 = 22